Help me to calculate the mean and range of the following clubs data:
Club A - 16.0,15.8,16.85,15.8,17.85,15.3,15.75,15.45
Club B - 15.95,16.9,16.85,17.85,15.25,17.5,17.0,17.25

Answer by rfadrogane(214)

Sol'n:
For Club A:
15.3
15.45
15.75
15.8
15.8
16
16.85
17.85
------
128.8
mean = 128.8/8
= 16.1 --ans.
range = 17.85-15.3
= 2.55 -- ans.
For Club B:
15.25
15.95
16.85
16.9
17
17.25
17.5
17.85
------
134.55
mean = 134.55/8
= 16.82 --ans.
range = 17.85-15.25
= 2.6 ---ans
thus, club B has higher range and mean